OpenStudy (anonymous):

solve system of equations: y^2= x x+2y+3= 0

OpenStudy (radar):

substituting y^2 for x in the second equation; y^2 +2y +3 = 0 solve for y using quadratic formula. then solve for x substituting the values obtained for y, in one of the original equation.

OpenStudy (anonymous):

yea I did that but I get imaginary numbers for y. so how do i get the x values ? i know i have to substitute them in the other equation but what should be the final answer ?

OpenStudy (anonymous):

i get y= -1 + 1i sqrt 2 and y = -1- 1i sqrt 2 for x i get x= -1 plus or minus 2i sqrt 2

OpenStudy (radar):